# Stellar-gas kinematic misalignments in eagle: lifetimes and longevity of misaligned galaxies

### v1

@@ Abstract
The dominant processes by which galaxies replenish their cold gas reservoirs remain disputed, especially in massive galaxies. Stellar-gas kinematic misalignments offer an opportunity to study these replenishment processes. However, observed distributions of these misalignments conflict with current models of gas replenishment in early-type galaxies (ETGs), with longer relaxation timescales suggested as a possible solution. We use the eagle simulation to explore the relaxation of unstable misaligned gas in galaxies with masses of $M_{*}\geqslant\mathrm{10^{9.5}}$ M between $10<z<$ 1. We extract misalignments from formation to relaxation providing a sample of $\sim 3200$ relaxations. We find relaxation timescales tend to be short-duration, with median lifetimes of $\sim 0.5$ Gyr, though with a notable population of unstable misalignments lasting $\gtrsim 1$ Gyr. Relaxation time distributions show a log-linear relationship, with $\approx 20\%$ of unstable misalignments persisting for $\gtrsim 3$ torquing times. Long-lived unstable misalignments are predominantly found in galaxies with higher stellar masses, lower star-forming gas fractions, higher ongoing gas inflow, and which reside in the centres of dense environments. Mergers only cause $\sim 17\%$ of unstable misalignments in eagle. We conclude that, at least in eagle, unstable kinematic misalignments are not predominantly driven by gas-rich minor mergers. Additionally, processes that significantly extend relaxation times are not dominant in the galaxy population. Instead, we see a diverse formation pathway for misalignments such as through hot halo cooling.
@@
pubyear: 2024pagerange: Stellar-gas kinematic misalignments in eagle: lifetimes and longevity of misaligned galaxiesA
@@ 1 Introduction
Stellar-gas kinematic misalignments (henceforth ‘misaligned galaxies’) have been one laboratory in which the origins of cold gas replenishment has been studied (e.g. Sarzi et al., 2006; Davis et al., 2011; Davis & Bureau, 2016; Bryant et al., 2019; Ruffa et al., 2019b; Duckworth et al., 2020a). Observationally, misalignments are found in $\approx 11\%$ of the galaxy population at $z\sim 0.1$ (Bryant et al., 2019) and show a strong morphological dependence, with $\sim 30-40\%$ of ETGs and $\sim 2-5\%$ of LTGs showing such features (e.g. Davis et al., 2011; Chen et al., 2016; Duckworth et al., 2019; Bryant et al., 2019; Ruffa et al., 2019a; Ristea et al., 2022; Raimundo et al., 2023). Cosmological hydrodynamical simulations, such as eagle (Schaye et al., 2015; Crain et al., 2015) and horizon-agn (Dubois et al., 2014), have also proven to be invaluable to understanding the physical drivers and timescales involved in misalignments, owing to their ability to trace galaxies over time. These have been able to recreate the observed misalignment populations (e.g. Duckworth et al., 2020a; Khim et al., 2020; Casanueva et al., 2022).
The existence of stellar-gas misalignments is predicted by our current Lambda Cold Dark Matter ($\Lambda$CDM) cosmological model. Stellar and gaseous components are expected to form an aligned discy structure as a natural consequence of galaxy formation (see tidal torque theory; TTT by Peebles, 1969; Doroshkevich, 1970; White, 1984). As the galaxy evolves, stellar and gas components may decouple and become misaligned, provided that the angular momentum of the existing gas disc is overcome and a new, misaligned disc can form. The stochastic nature of external gas accretion and ability, to first order, accrete with a random angular momentum orientation has led external accretion to become the favoured dominant driver behind many misalignments (e.g. Davis et al., 2011; van de Voort et al., 2015; Starkenburg et al., 2019; Duckworth et al., 2020b; Khim et al., 2021; Casanueva et al., 2022).
@@
A key area of contention lies between the observed shape of the misalignment fraction distributions and those predicted by analytical models. As explored in detail in Davis & Bureau (2016), assuming mergers are the primary driver of misalignments alongside theoretical relaxation timescales (in the absence of smooth accretion), current models significantly overpredict the number of relaxed systems. Recent theoretical findings point to a significantly more diverse formation path for misaligned systems such as through galaxy fly-bys and halo cooling onto gas-poor systems (e.g. Lagos et al., 2015; Bryant et al., 2019; Khim et al., 2021). Likewise, the inclusion of smooth accretion onto an existing misaligned gas disc has been shown to significantly enhance the misalignment timescales beyond theoretical timescales (e.g. van de Voort et al., 2015; Lagos et al., 2015; Ruffa et al., 2019b). This misaligned smooth accretion may be supplied through halo cooling (e.g. Negri et al., 2014; Lagos et al., 2015; Ruffa et al., 2019b; Duckworth et al., 2020a), provided the galaxy hosts a gaseous halo that is kinematically misaligned with respect to its stellar component (e.g. Padilla et al., 2014; Duckworth et al., 2020a). Misaligned smooth accretion may also be supplied from tidal tails from a recent merger which can significantly extend the longevity of a misalignment beyond $>2$ Gyr (van de Voort et al., 2015). Taken together, these results suggest the processes by which galaxies replenish their cold gas may be more diverse than previously thought. Thus, by investigating the timescales of misalignments in hydrodynamical simulations we can gain insights into gas replenishment processes that may be challenging to detect observationally.
@@
Thus far, the exploration of misalignment timescales has been largely limited to small numbers of galaxies with idealised behaviour. One notable exception is from Khim et al. (2021) who used the horizon-agn simulation to study the lifetimes of misalignments in a sizeable galaxy population. Using a sample of $\sim 2200$ misaligned galaxies at $z=0.52$ and measuring the population decay, they find a significant population of galaxies with misalignment lifetimes of $>2$ Gyr with a strong morphological dependence. ETGs were found with characteristic decay timescales of $\approx 2.3$ Gyr compared to $\approx 0.8$ Gyr for LTGs. Overall, galaxies with lower gas fractions, more early-type morphologies, and those residing in denser environments were found to have longer decay timescales. As these properties are common among ETGs (Dressler, 1980), this is thought to amplify the higher misalignment fractions among ETG populations. Additionally, misalignments formed from interactions between the galaxy and a dense medium (e.g. ram-pressure stripping; Gunn & Gott, 1972) were found to be the most long-lived, with those caused by mergers, galaxy interactions, and remaining processes including halo cooling and filament accretion resulting in faster relaxations.
@@
Another noticeable exception is from Casanueva et al. (2022) who used the eagle simulation to investigate the origin of misaligned galaxies for a representative galaxy population. They find galaxies that remain misaligned between two neighbouring simulation snapshots are typically more spheroidal, dispersion-dominated, and gas-poor. Misalignment formation was found to correlate most strongly with a reduction in cold gas fraction, which was attributed to gas stripping from nearby galaxy interactions. Mergers were found to coincide with misalignment formation in only $\approx 4\%$. Alongside comparable external accretion rates in aligned galaxies, Casanueva et al. (2022) suggest external processes are not the dominant driver of misalignments in eagle, in agreement with Khim et al. (2021).
@@
In this paper, we use the eagle (Evolution and Assembly of GaLaxies and their Environments) suite of cosmological hydrodynamical simulations (Schaye et al., 2015; Crain et al., 2015) to investigate the persistence of stellar-gas kinematic misalignments for a representative galaxy population from $10<z<$ 1. Importantly, our approach differs from that of Khim et al. (2021) as we extract a population of misalignments over a complete window from formation to relaxation, rather than characterising relaxation timescales using a misalignment population decay from a given redshift. Additionally, we study these misalignments as 3D angles, rather than in projection, and treat stable co- and counter-rotating regimes separately. In this way, we create a reliable estimate of the misalignment timescales for a large galaxy sample for the first time. Specifically, we aim to answer the following questions: 1) whether significantly longer relaxation timescales (of order $\sim 20-80$ dynamical timescales) as suggested by Davis & Bureau (2016) are seen within eagle; 2) the degree to which relaxation times vary with galaxy properties and environment; and 3) to what extent are mergers associated with misalignment formation.
@@ 2 Theoretical relaxation timescales
As outlined by Tohline et al. (1982), a misaligned disc of gas can be modelled by a series of rings embedded within an ellipsoidal gravitational potential well. These rings of gas have a circular velocity $V_{\mathrm{rot}}$ at a radius $R$. The non-zero angular momentum component of this gas will experience a net torque from the existing mass distribution. This results in a precession rate given by {ceqn}
@@
where $\epsilon$ is the stellar ellipticity and the dynamical time of the galaxy is defined as $t_{\mathrm{dyn}}=2\pi R/V_{\mathrm{rot}}$. For a typical ETG with $\epsilon\approx 0.2$, a disc size of $R\approx 1$ kpc, and $V_{\mathrm{rot}}\approx 200$ km s-1, this implies $t_{\mathrm{relax}}\approx 5t_{\mathrm{dyn}}$, corresponding to $\sim 0.2$ Gyr.
@@ 3 The eagle simulation
The Evolution and Assembly of GaLaxies and their Environments [^1] A full description of models and parameters can be found in Schaye et al. (2015) and Crain et al. (2015), with publicly available data products available from http://eagle.strw.leidenuniv.nl and http://www.eaglesim.org/ and described in McAlpine et al. (2017) (eagle) is a suite of cosmological hydrodynamical simulations described in Schaye et al. (2015) and Crain et al. (2015). These simulations use a modified version of the parallel N-body smoothed particle hydrodynamics (SPH) code Gadget-3 (Springel et al., 2005, 2008) known as Anarchy (see Schaller et al. 2015 for description and analysis). The initial conditions for these simulations were generated using Panphasia (Jenkins & Booth, 2013). A $\Lambda$CDM framework is also assumed, with $\Omega_{\rm{m}}=0.307$, $\Omega_{\rm{\Lambda}}=0.693$, $\Omega_{\rm{b}}=0.04825$, $h=0.6777$, $\sigma_{\rm{8}}=0.8288$, $n_{\rm{s}}=0.9611$ and $Y=0.248$ (Planck Collaboration et al., 2014).
eagle uses sub-grid models to capture physical processes below the resolution limit ($\sim 0.7$ pkpc) of the simulation. These include radiative cooling and photoheating (Wiersma et al., 2009a), star formation (Schaye & Dalla Vecchia, 2008), stellar evolution and interstellar medium (ISM) enrichment (Wiersma et al., 2009b), the stochastic thermal stellar feedback by Dalla Vecchia & Schaye (2012), and black hole growth through accretion and mergers and feedback (Springel et al., 2005; Rosas-Guevara et al., 2015). The eagle sub-grid parameters are tuned to reproduce observations of the galaxy stellar mass function and galaxy sizes relation at $z\sim 0$. The eagle simulations have been successful in reproducing observational mass, size, and morphology distributions (e.g., Schaye et al., 2015; Correa et al., 2017; Trayford & Schaye, 2019; Hill et al., 2021).
Galaxy evolution can be traced using the subhalo merger trees by Qu et al. (2017) that were generated using the D-Trees algorithm (Jiang et al., 2014). The latter links subhalos by establishing the location of the majority of all particles assigned to a particular subhalo between consecutive snapshots [^2] For a more detailed explanation of the D-Trees algorithm and merger tree construction see Jiang et al. (2014) and Qu et al. (2017), respectively. Mergers are treated hierarchically such that two (or more) progenitor galaxies undergoing a merger will form a single descendant galaxy. A merger is defined to occur once the D-Trees algorithm can no longer identify a progenitor subhalo among the subfind catalogues. At this point the subfind algorithm can no longer distinguish the progenitor density profiles, at which point particles from the secondary galaxy are reassigned to the primary galaxy. Occasionally during mergers, the algorithm may switch the primary and secondary galaxies if these are in very close proximity (‘Subhalo switching problem’; see Qu et al., 2017, for details). Regarding this work, a switch between coalescing galaxies in which one galaxy is misaligned may distort the measured misalignment duration. To minimise this effect, we exclude galaxies that experience a spontaneous stellar mass decrease of $>50\%$ whilst misaligned. This ensures a spontaneous switch to a low-mass merging satellite is not falsely included as a misalignment. However, we do not find this choice to meaningfully affect our sample or results.
Table 1: Parameters of the Ref-L100N1504 simulation from the eagle suite. From top to bottom, the rows report: co-moving box size; initial dark matter (DM) particle count (initially there is an equal number of baryonic particles); initial particle mass of gas particles; initial particle mass of DM particles; the Hubble parameter ($h=H_{0}/100$ km s-1 Mpc-1); co-moving, Plummer-equivalent gravitational softening length; maximum physical gravitational softening length.
@@
total particle count ×$\times 1504^{3}$
@@
The eagle simulation is stored in 400 intervals between $z\sim 20$ to $z=0$. Of these, 29 are "snapshots" with time intervals from $\approx 0.3$ to 1 Gyr, which record complete particle properties alongside extensive halo and subhalo property catalogues. Additionally, 400 higher time-cadence "snipshots" are stored with limited catalogues of particle and subhalo properties, though with enough information to calculate misalignment angles. We use 200 of these 400 snipshots for which merger trees were run by Crain et al. (2017), with a mean time cadence of $\approx 0.12$ Gyr within our target redshift range. These allow us to distinguish between faster ($\sim 0.1$ Gyr) and longer-lived ($\gtrsim 1$ Gyr) misalignments. As we wish to focus on the period in which galaxy growth occurs largely linearly, we limit our data analysis to $10<z<$ 1 which provides a wide temporal window over which to measure relaxation timescale distributions. For this work we make use of the largest reference simulation, Ref-L100N1504, which has a box size of 100 co-moving Mpc3 (henceforth ckpc for co-moving and pkpc for proper length-scales). Parameters specific to the Ref-L100N1504 are summarised in Table 1.
Figure 1: Galaxy stellar mass function at $z=0.1$ of the total galaxy population in the Ref-L100N1504 simulation (solid purple line) in comparison with our sub-sample of galaxies (dashed red line), which were selected using the criteria listed in Section 4. There is no mass bias in our sub-sample compared to the general galaxy population.

### v2

@@ Abstract
The dominant processes by which galaxies replenish their cold gas reservoirs remain disputed, especially in massive galaxies. Stellar-gas kinematic misalignments offer an opportunity to study these replenishment processes. However, observed distributions of these misalignments conflict with current models of gas replenishment in early-type galaxies (ETGs), with longer relaxation timescales suggested as a possible solution. We use the eagle simulation to explore the relaxation of unstable misaligned gas in galaxies with masses of $M_{*}\geqslant\mathrm{10^{9.5}}$ M between $0<z<1$. We extract misalignments from formation to relaxation providing a sample of $\sim 3200$ relaxations. We find relaxation timescales tend to be short-duration, with median lifetimes of $\sim 0.5$ Gyr, though with a notable population of unstable misalignments lasting $\gtrsim 1$ Gyr. Relaxation time distributions show a log-linear relationship, with $\approx 20$ per cent of unstable misalignments persisting for $\gtrsim 3$ torquing times. Long-lived unstable misalignments are predominantly found in galaxies with higher stellar masses, lower star-forming gas fractions, higher ongoing gas inflow, and which reside in the centres of dense environments. Mergers only cause $\approx 10$ per cent of unstable misalignments among galaxies at $z<0.35$, and $\approx 21$ per cent at $0.35<z<1.0$ in eagle. We conclude that, at least in eagle, unstable kinematic misalignments are not predominantly driven by gas-rich minor mergers at any redshift probed. Additionally, processes that significantly extend relaxation times are not dominant in the galaxy population. Instead, we see a diverse formation pathway for misalignments such as through hot halo cooling.
@@
pubyear: 2025pagerange: Stellar-gas kinematic misalignments in eagle: lifetimes and longevity of misaligned galaxiesA
@@ 1 Introduction
Stellar-gas kinematic misalignments (henceforth ‘misaligned galaxies’) have been one laboratory in which the origins of cold gas replenishment has been studied (e.g. Sarzi et al., 2006; Davis et al., 2011; Davis & Bureau, 2016; Bryant et al., 2019; Ruffa et al., 2019b; Duckworth et al., 2020a). Observationally, misalignments are found in $\approx 11$ per cent of the galaxy population at $z\sim 0.1$ (Bryant et al., 2019) and show a strong morphological dependence, with $\sim 30-40$ per cent of ETGs and $\sim 2-5$ per cent of LTGs showing such features (e.g. Davis et al., 2011; Chen et al., 2016; Duckworth et al., 2019; Bryant et al., 2019; Ruffa et al., 2019a; Ristea et al., 2022; Raimundo et al., 2023). Cosmological hydrodynamical simulations, such as eagle (Schaye et al., 2015; Crain et al., 2015) and horizon-agn (Dubois et al., 2014), have also proven to be invaluable to understanding the physical drivers and timescales involved in misalignments, owing to their ability to trace galaxies over time. These have been able to recreate the observed misalignment populations (e.g. Duckworth et al., 2020a; Khim et al., 2020; Casanueva et al., 2022).
[…]
@@
A key area of contention lies between the observed shape of the misalignment fraction distributions and those predicted by analytical models. As explored in detail in Davis & Bureau (2016), assuming mergers are the primary driver of misalignments alongside theoretical relaxation timescales (in the absence of smooth accretion), current models significantly overpredict the number of relaxed systems. Recent theoretical findings point to a significantly more diverse formation path for misaligned systems. These include the accretion of gas stripped from a nearby low-mass satellite, perturbations of the existing co-rotating gas induced from fly-bys or ram-pressure stripping, and halo cooling onto gas-poor systems (e.g. Lagos et al., 2015; Bryant et al., 2019; Khim et al., 2021; Cenci et al., 2024). Likewise, the inclusion of smooth accretion onto an existing misaligned gas disc has been shown to significantly enhance the misalignment timescales beyond theoretical timescales (e.g. van de Voort et al., 2015; Lagos et al., 2015; Ruffa et al., 2019b). This misaligned smooth accretion may be supplied through halo cooling (e.g. Negri et al., 2014; Lagos et al., 2015; Ruffa et al., 2019b; Duckworth et al., 2020a), provided the galaxy hosts a gaseous halo that is kinematically misaligned with respect to its stellar component (e.g. Padilla et al., 2014; Duckworth et al., 2020a). Misaligned smooth accretion may also be supplied from tidal tails from a recent merger which can significantly extend the longevity of a misalignment beyond >2>2 Gyr (van de Voort et al., 2015). Taken together, these results suggest the processes by which galaxies replenish their cold gas may be more diverse than previously thought. Thus, by investigating the timescales of misalignments in hydrodynamical simulations we can gain insights into gas replenishment processes that may be challenging to detect observationally.
@@
Thus far, the exploration of misalignment timescales has been largely limited to small numbers of galaxies with idealised behaviour. One notable exception is from Khim et al. (2021) who used the horizon-agn simulation to study the lifetimes of misalignments in a sizeable galaxy population. Using a sample of $\sim 2200$ misaligned galaxies at $z=0.52$ and measuring the population decay, they find a significant population of galaxies with misalignment lifetimes of >2>2 Gyr with a strong morphological dependence. ETGs were found with characteristic decay timescales of $\approx 2.3$ Gyr compared to $\approx 0.8$ Gyr for LTGs. Overall, galaxies with lower gas fractions, more early-type morphologies, and those residing in denser environments were found to have longer decay timescales. As these properties are common among ETGs (Dressler, 1980), this is thought to amplify the higher misalignment fractions among ETG populations. Additionally, misalignments formed from interactions between the galaxy and a dense medium (e.g. ram-pressure stripping; Gunn & Gott, 1972) were found to be the most long-lived, with those caused by mergers, galaxy interactions, and remaining processes including halo cooling and filament accretion resulting in faster relaxations.
@@
Another noticeable exception is from Casanueva et al. (2022) who used the eagle simulation to investigate the origin of misaligned galaxies for a representative galaxy population. They find galaxies that remain misaligned between two neighbouring simulation snapshots are typically more spheroidal, dispersion-dominated, and gas-poor. Misalignment formation was found to correlate most strongly with a reduction in cold gas fraction, which was attributed to gas stripping from nearby galaxy interactions. Mergers were found to coincide with misalignment formation in only $\approx 4$ per cent. Alongside comparable external accretion rates in aligned galaxies, Casanueva et al. (2022) suggest external processes are not the dominant driver of misalignments in eagle, in agreement with Khim et al. (2021).
@@
In this paper, we use the eagle (Evolution and Assembly of GaLaxies and their Environments) suite of cosmological hydrodynamical simulations (Schaye et al., 2015; Crain et al., 2015) to investigate the persistence of stellar-gas kinematic misalignments for a representative galaxy population from $0<z<1$. Importantly, our approach differs from that of Khim et al. (2021) as we extract a population of misalignments over a complete window from formation to relaxation, rather than characterising relaxation timescales using a misalignment population decay from a given redshift. Additionally, we study these misalignments as 3D angles, rather than in projection, and treat stable co- and counter-rotating regimes separately. In this way, we create a reliable estimate of the misalignment timescales for a large galaxy sample for the first time. Specifically, we aim to answer the following questions: 1) whether significantly longer relaxation timescales (of order $\sim 20-80$ dynamical timescales) as suggested by Davis & Bureau (2016) are seen within eagle; 2) the degree to which relaxation times vary with galaxy properties and environment; and 3) to what extent are mergers associated with misalignment formation.
@@ 2 Theoretical relaxation timescales
As outlined by Tohline et al. (1982), a misaligned disc of gas can be modelled by a series of rings embedded within an ellipsoidal gravitational potential well. These rings of gas have a circular velocity $V_{\mathrm{rot}}$ at a radius $R$. The non-zero angular momentum component of this gas will experience a net torque from the existing mass distribution. This results in a precession rate given by \@mathmargin
@@
where $\epsilon$ is the stellar ellipticity and the dynamical time of the galaxy is defined as $t_{\mathrm{dyn}}=2\pi R/V_{\mathrm{rot}}$. For a typical ETG with $\epsilon\approx 0.2$, a disc size of $R\approx 1$ kpc, and $V_{\mathrm{rot}}\approx 200$ km s-1, this implies $t_{\mathrm{relax}}\approx 5t_{\mathrm{dyn}}$, corresponding to $\sim 0.2$ Gyr. This highlights the expected short-lived lifetimes of stellar-gas kinematic misalignments.
@@ 3 The eagle simulation
[…]
eagle uses sub-grid models to capture physical processes below the resolution limit ($\sim 0.7$ pkpc) of the simulation. These include radiative cooling and photoheating (Wiersma et al., 2009a), star formation (Schaye & Dalla Vecchia, 2008), stellar evolution and interstellar medium (ISM) enrichment (Wiersma et al., 2009b), the stochastic thermal stellar feedback by Dalla Vecchia & Schaye (2012), and black hole growth through accretion and mergers and feedback (Springel et al., 2005; Rosas-Guevara et al., 2015). The eagle sub-grid parameters are tuned to reproduce observations of the galaxy stellar mass function and galaxy sizes relation at $z\sim 0$. The eagle simulations have been successful in reproducing observational mass, size, and morphology distributions (e.g., Schaye et al., 2015; Correa et al., 2017; Trayford & Schaye, 2019; Pfeffer et al., 2023).
Table 1: Parameters of the Ref-L100N1504 simulation from the eagle suite. From top to bottom, the rows report: co-moving box size; initial dark matter (DM) particle count (initially there is an equal number of baryonic particles); initial particle mass of gas particles; initial particle mass of DM particles; co-moving, Plummer-equivalent gravitational softening length; maximum physical gravitational softening length.
@@
total particle count $\times 1504^{3}$
@@
Galaxy evolution can be traced using the subhalo merger trees by Qu et al. (2017) that were generated using the D-Trees algorithm (Jiang et al., 2014). The latter links subhalos by establishing the location of the majority of all particles assigned to a particular subhalo between consecutive snapshots [^2] For a more detailed explanation of the D-Trees algorithm and merger tree construction see Jiang et al. (2014) and Qu et al. (2017), respectively. Mergers are treated hierarchically such that two (or more) progenitor galaxies undergoing a merger will form a single descendant galaxy. A merger is defined to occur once the D-Trees algorithm can no longer identify a progenitor subhalo among the subfind catalogues. At this point the subfind algorithm can no longer distinguish the progenitor density profiles, at which point particles from the secondary galaxy are reassigned to the primary galaxy. Occasionally during mergers, the algorithm may switch the primary and secondary galaxies if these are in very close proximity (‘Subhalo switching problem’; see Qu et al., 2017, for details). Regarding this work, a switch between coalescing galaxies in which one galaxy is misaligned may distort the measured misalignment duration. To minimise this effect, we exclude galaxies that experience a spontaneous stellar mass decrease of >50>50 per cent whilst misaligned. This ensures a spontaneous switch to a low-mass merging satellite is not falsely included as a misalignment. However, we do not find this choice to meaningfully affect our sample or results.
The eagle simulation is stored in 400 intervals between $z\sim 20$ to $z=0$. Of these, 29 are "snapshots" with time intervals from $\approx 0.3$ to 1 Gyr, which record complete particle properties alongside extensive halo and subhalo property catalogues. Additionally, 400 higher time-cadence "snipshots" are stored with limited catalogues of particle and subhalo properties, though with enough information to calculate misalignment angles. We use 200 of these 400 snipshots for which merger trees were run by Crain et al. (2017), with a mean time cadence of $\approx 0.12$ Gyr within our target redshift range. These allow us to distinguish between faster ($\sim 0.1$ Gyr) and longer-lived ($\gtrsim 1$ Gyr) misalignments. As we wish to focus on the period in which galaxy growth occurs largely linearly, we limit our data analysis to $0<z<1$ which provides a wide temporal window over which to measure relaxation timescale distributions. For this work we make use of the largest reference simulation, Ref-L100N1504, which has a box size of 100 co-moving Mpc3 (henceforth ckpc for co-moving and pkpc for proper length-scales). Parameters specific to the Ref-L100N1504 are summarised in Table 1.
Figure 1: Galaxy stellar mass function at $z=0.1$ of the total galaxy population (solid purple line) in the Ref-L100N1504 simulation in comparison with our sub-sample of galaxies (dashed red line) selected using the criteria listed in Section 4. There is no mass bias in our sub-sample compared to the general galaxy population.